# ChemKANs for Combustion Chemistry Modeling and Acceleration

### v1

@@ 3.3 Current limitations and areas for further research
While we have demonstrated that our ChemKANs provide remarkable inference capability and expressivity, there remain drawbacks. The total speedup reported in Table 1 is ultimately not as large as we believe that it could be. While still more than competitive with that of ChemNODE especially considering the complete thermochemical source term with low-concentration radicals that ChemKAN provides, it appears underwhelming in light of the rest of the significant performance gains enabled by ChemKAN in other comparisons throughout this work. That being said, we believe that the reported speedup is a conservative lower bound on the potential for ChemKAN and KANs in general. The original KANs [23] are at the time of original submission of the current work less than a year old and are known in the literature to be much slower than comparable MLPs (this is even acknowledged by the authors of the original KAN paper [23]). However, extensive research is actively ongoing to resolve this issue including parameter efficiency improvements [33, 51, 35] and prediction acceleration techniques [38, 52, 53, 54, 34, 39, 55]. It’s unclear which methods will ultimately prevail, but the recent emergence of KANs and the large amount of work proposing various techniques for their acceleration suggest significant promise in the near future for substantial and relatively lower-effort acceleration, compared to the more mature and well-developed MLPs where we believe it reasonable to expect a slower pace of future development.
@@
We have additionally in this section compared a baseline ChemKAN implementation against a baseline ChemNODE implementation. Later works exist that appear to successfully combine ChemNODE with augmented loss functions, autoencoders, and latent space time stepping. The most recent, which includes all of these techniques, is reported in Ref. [42] (“Phy-ChemNODE”). We do not draw comparisons between the current ChemKAN implementation and the larger-scale, combined-methodology results in [42], as our current aim is to compare the pure performance of the ChemKAN against the MLPs that underlie both ChemNODEs and Phy-ChemNODEs. All further augmentations carried out in Phy-ChemNODE that go beyond this baseline can be replicated with ChemKANs, and an interesting target of future studies may be on quantifying the performance gains of ChemKAN when applied in tandem with other advanced neural network structures.
4 Conclusions
@@
This work introduced ChemKANs, a novel physics-informed machine learning technique based on the general KAN-ODE framework with a specialized structure tailored explicitly for chemical kinetic modeling and acceleration. Its two-part design allows for application in isothermal or exothermic systems through a kinetic core and an optional thermodynamic superstructure that can be trained and applied as a single cohesive framework. Model inference in a preliminary biodiesel synthesis case using only the kinetic core revealed that ChemKANs are remarkably robust to overfitting. Fair comparisons against a generic Deep Operator Network approach revealed that with increasingly bulky parameterizations, ChemKANs retained plateaud optimal loss metrics, while DeepONets saw minor further decreases in training loss accompanied with an increase in testing loss, as might be expected from a standard deep learning approach. With added noise, this difference was further exacerbated, with the DeepONet achieving low training loss through jagged fits to the noise itself, while the ChemKAN converged its training, testing, and noise-free testing performance with smooth fits to the hidden underlying data. While alone promising, these results are of particular interest given prior works in the literature that have casted doubts on the effectiveness of KAN structures on noisy functions.
@@
In a second case, ChemKANs were demonstrated as efficient acceleration surrogates for learning chemical source terms in a hydrogen combustion case. A two-stage training process for the kinetic core and thermodynamic superstructure enabled a single, 344-parameter ChemKAN to accurately learn complete solution profiles across a range of hydrogen-air homogeneous reactor initial conditions, a significant reduction in parameter and network bulk compared to previous MLP-based neural ODE approaches that required 637 parameters to learn a truncated set of solution profiles. Timing comparisons against the detailed mechanism revealed a 2 speedup when using the ChemKAN surrogate model, which is significant for downstream applications of the hydrogen combustion surrogate learned here (for example, 3-D turbulent reacting flow). In summary, we find that ChemKANs are a promising tool for both dynamical system modeling and acceleration tasks in combustion chemistry. In doing so, we have also successfully advanced the underlying KAN-ODE framework to much larger, practical systems than have been studied previously. We hope that these promising preliminary case studies motivate future implementation of KAN layers and modules in combustion machine learning applicatins.
Acknowledgement
@@
CRediT authorship contribution statement
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ III.3 Current limitations and areas for further research
While we have demonstrated that our ChemKANs provide remarkable inference capability and expressivity, there remain drawbacks. The total speedup reported in Table 1 is ultimately not as large as we believe that it could be. While still more than competitive with that of ChemNODE, especially considering the complete thermochemical source term with low-concentration radicals that ChemKAN provides, it appears underwhelming in light of the rest of the significant performance gains enabled by ChemKAN in other comparisons throughout this work. That being said, we believe that the reported speedup is a conservative lower bound on the potential for ChemKAN and KANs in general. The original KANs [23] are, at the time of original submission of the current work, barely a year old and are known in the literature to be much slower than comparable MLPs (this is even acknowledged by the authors of the original KAN paper [23]). However, extensive research is actively ongoing to resolve this issue, including parameter efficiency improvements [35, 53, 37] and prediction acceleration techniques [40, 54, 55, 56, 36, 41, 57]. It is unclear which methods will ultimately prevail, but the recent emergence of KANs and the large amount of work proposing various techniques for their acceleration suggest significant promise in the near future for substantial and relatively lower-effort acceleration, compared to the more mature and well-developed MLPs where we believe it reasonable to expect a slower pace of future development.
@@
We have additionally in this section compared a baseline ChemKAN implementation against a baseline ChemNODE implementation. Later works exist that appear to successfully combine ChemNODE with augmented loss functions, autoencoders, and latent space time stepping. The most recent, “Phy-ChemNODE”, includes all of these techniques [44]. We do not draw comparisons between the current ChemKAN implementation and the larger-scale, combined-methodology results reported there, as our current aim is to compare the pure performance of ChemKAN against the MLPs that underlie both ChemNODEs and Phy-ChemNODEs. All further augmentations carried out in Phy-ChemNODE that go beyond this baseline can be replicated with ChemKANs, and an interesting target of future studies may be to quantify the performance gains of ChemKAN when applied in tandem with other advanced neural network structures.
Conclusions
@@
This work introduced ChemKANs, a novel physics-informed machine learning technique based on the general KAN-ODE framework with a specialized structure tailored explicitly for chemical kinetic modeling and acceleration. Its two-part design allows for application in isothermal or exothermic systems through a kinetic core and an optional thermodynamic superstructure that can be trained and applied as a single cohesive framework. Model inference in a preliminary biodiesel synthesis case using only the kinetic core revealed that ChemKANs are remarkably robust to overfitting. Fair comparisons against a generic DeepONet approach revealed that with increasingly bulky parameterizations, ChemKANs retained plateaued optimal loss metrics, while DeepONets saw minor further decreases in training loss accompanied by an increase in testing loss, as might be expected from a standard deep learning approach. With added noise, this difference was further exacerbated, with the DeepONet achieving low training loss through jagged fits to the noise itself, while the ChemKAN converged its training, testing, and noise-free testing performance with smooth fits to the hidden underlying data. While alone promising, these results are of particular interest given prior works in the literature that have cast doubts on the effectiveness of KAN structures on noisy functions.
@@
In a second case, ChemKANs were demonstrated as efficient acceleration surrogates for learning chemical source terms in a hydrogen combustion case. A two-stage training process for the kinetic core and thermodynamic superstructure enabled a single, 344-parameter ChemKAN to accurately learn complete solution profiles across a range of hydrogen-air homogeneous reactor initial conditions, a significant reduction in parameter and network bulk compared to previous MLP-based neural ODE approaches that required 637 parameters to learn a truncated set of solution profiles. Timing comparisons against the detailed mechanism revealed a 2 speedup when using the ChemKAN surrogate model, which is significant for downstream applications of the hydrogen combustion surrogate learned here (for example, 3-D turbulent reacting flow). In summary, we find that ChemKANs are a promising tool for both dynamical system modeling and acceleration tasks in combustion chemistry. In doing so, we have also successfully advanced the underlying KAN-ODE framework to much larger, practical systems than had been studied previously. We hope that these promising preliminary case studies motivate future implementation of ChemKAN layers and modules in combustion and chemical kinetic machine learning applications.
Author contributions
Conflicts of interest
There are no conflicts to declare.
Data Availability
The data used to produce the results in this manuscript are available on request.
Acknowledgements